### v1

# DiffusionAvatars: Deferred Diffusion for High-fidelity 3D Head Avatars

Abstract
DiffusionAvatars synthesizes a high-fidelity 3D head avatar of a person, offering intuitive control over both pose and expression. We propose a diffusion-based neural renderer that leverages generic 2D priors to produce compelling images of faces. For coarse guidance of the expression and head pose, we render a neural parametric head model (NPHM) from the target viewpoint, which acts as a proxy geometry of the person. Additionally, to enhance the modeling of intricate facial expressions, we condition DiffusionAvatars directly on the expression codes obtained from NPHM via cross-attention. Finally, to synthesize consistent surface details across different viewpoints and expressions, we rig learnable spatial features to the head’s surface via TriPlane lookup in NPHM’s canonical space.
We train DiffusionAvatars on RGB videos and corresponding tracked NPHM meshes of a person and test the obtained avatars in both self-reenactment and animation scenarios. Our experiments demonstrate that DiffusionAvatars generates temporally consistent and visually appealing videos for novel poses and expressions of a person, outperforming existing approaches.
(a) Avatar creation
(b) Avatar animation
Figure 1: Given a set of multi-view videos and corresponding tracked meshes, we build a DiffusionAvatar of a person. Our method translates expressions of a morphable model into realistic facial appearances of a person while also providing control over the viewpoint.
1 Introduction
A significant problem in computer vision and graphics is the creation of photorealistic animatable avatars. Ideally, these avatars permit consistent video renderings while providing free control over desired expression, pose, and viewpoint. A digital 3D head avatar may then be used in various scenarios such as VR/AR applications, immersive teleconferencing, gaming, movie animation, language learning, or as a virtual assistant.
Unfortunately, creating a 3D avatar from which one can render photorealistic images from arbitrary viewpoints is challenging. The task becomes even more difficult when the avatar has to simultaneously offer precise control over poses and expressions, for instance given by a driving sequence from a different video or controlled manually. Technically, this leads to a 4D photometric reconstruction problem which is typically underconstrained. This presents animation systems with a challenging task that traditionally involves extensive efforts by 3D artists.
Many existing methods have tackled the challenge of facial image synthesis via 2D neural networks [29, 69, 24, 54, 66, 48]. In particular, Diffusion Models have recently been demonstrated to produce visually appealing imagery [21, 55, 57, 84, 81]. 2D models are great at generating photo-realistic images but generally do not offer the degree of control and temporal consistency needed for a high-quality 3D head avatar. On the other side, methods that reconstruct an animatable 3D representation of the head [90, 19, 88, 87] provide much better consistency but in many cases the resulting renderings do not have the same photo-realism as 2D models.
To this end, we introduce DiffusionAvatars, a novel approach for 3D head avatar creation, by combining the strong image synthesis capabilities of 2D diffusion models with the view consistency of a detailed 3D head representation. As a controllable 3D representation, we leverage the recent neural parametric head model (NPHM) [17] due to its detailed prediction of human head geometry, which can give us better geometric cues than established mesh-based 3DMMs [35, 51]. Inspired by neural textures [68], we further base face synthesis on person-specific learnable features rigged to the model’s surface via NPHM’s canonical space. Such surface-mapped features can effectively compensate for imperfect geometry and improve consistency across views by giving the neural renderer cues about corresponding surface points. To avoid learning a neural renderer from scratch, we base our architecture on a pre-trained latent diffusion model (LDM) [55] and convert it into an image-to-image translation model by following the ControlNet [84] paradigm to condition on the rasterized images of NPHM meshes. That way, we not only inherit a powerful image synthesis backbone but also benefit from the learned facial prior obtained from large 2D image datasets, which facilitates generalization to unseen expressions. Finally, we insert cross-attention blocks into the pre-trained LDM to condition the diffusion-based neural renderer on the expression codes of NPHM directly. This allows our model to produce more expressive renderings. Intuitively, the direct conditioning helps the diffusion model to distinguish subtle expression details while the rasterized NPHM renderings encode the viewpoint and the overall shape of the head. In summary, our contributions are as follows:
  • We present DiffusionAvatars, a diffusion-based neural renderer that leverages ControlNet to create animatable 3D head avatars.
  • We design a method for rigging learnable spatial features to the surface of the underlying NPHM via TriPlanes.
  • We propose direct expression conditioning via cross-attention to transfer detailed expressions from NPHM to the synthesized 3D head avatar.
2 Related Work
2.1 3D Face Animation
It is natural to approach the task of 3D Face Animation with an actual 3D representation. In the seminal work by [4], a 3D morphable model (3DMM) was introduced, which paved the way towards reconstructing animatable 3D head geometry from images or depth observations. More recent 3DMMs [35, 51, 70, 77, 7] still build upon this paradigm and have since been the drivers of face animation oriented applications with most methods leveraging a 3DMM to gain control over the expression and pose of the avatar. For example, Neural Head Avatars [19] and ROME [30] finetune the mesh topology of FLAME [35] to obtain more faithful and realistic mesh-based avatars. NeRFace [15], INSTA [90] and RigNeRF [1] build a radiance field that is controlled by a 3DMM. Other 3D representations such as points [88] or signed distance functions [87, 37] have also been explored for avatar creation.
In our work, we do not aim to learn a complete 3D representation. Instead, we utilize the 3D prior of the recent neural parametric head model (NPHM) [17] with a 2D rendering network for high-quality image synthesis.
2.2 Face Synthesis in 2D
Ever since the impactful Pix2Pix [89] work, it has been a popular approach to obtain control over synthesized images via 2D networks. In the face synthesis domain, methods based on Generative Adversarial Networks (GANs) [18] have received a lot of attention [29, 69, 24, 54, 66, 48]. To obtain better control over head pose and camera viewpoint, a common technique is to combine powerful 2D image synthesis networks with a controllable 3D head proxy [8, 63, 65, 2, 45, 10, 75, 68, 36, 67]. Another line of work aims at directly generating consistent videos by video-to-video translation [31, 72] or even video stylization [78, 26, 14]. Most related to our method is Deferred Neural Rendering [68] which allows facial animation by learning a neural renderer in order to decode neural textures rigged to Basel Face Model [51]. While 2D-based approaches produce visually appealing frames, they often struggle with synthesizing consistent images across both views and time or provide only limited viewpoint control.
Our method aims to improve consistency by utilizing a more powerful implicit 3D representation as geometric proxy and provides better expression generalization by utilizing a diffusion model as a synthesis backbone.
2.3 Controllable 2D Diffusion
Recently, Diffusion models emerged as powerful 2D image generation models. While they demonstrate superior capabilities in generating high-fidelity and diverse 2D content from text prompts [21, 55, 57], it has been a key challenge to leverage their power for other tasks. Several works extend 2D diffusion models to text-guided video generation by employing a multi-stage approach [5, 22, 73, 60], generating frames autoregressively [23] or stylizing an input video [74, 16, 79, 76]. In the human domain, diffusion models have already been applied to face editing [25], animation [81, 62] and bodies [64]. Other works, such as ControlNet [84], IPAdapter [80] or T2I-Adapter [46], allow fine-tuning a pre-trained diffusion model on additional controls such as landmarks or depth cues, opening a wide space of possible applications for 2D face editing. Most similar to our method, DiffusionRig [11] proposes a diffusion model conditioned on rasterized FLAME meshes to provide 3D animation control over an avatar created from a personal photo album. While the synthesized images or videos of these approaches already show great controllability and visual quality, they typically lack consistency due to the absence of an underlying accurate 3D representation.
2.4 Diffusion for 3D Faces
Following a different approach, several works explore diffusion for generic 3D generation [52, 40, 47]. This has also been extended to 3D head generation [71, 20, 83], text-guided 3D head editing [49] or 3D body generation [82]. Some works also generate animatable 3D asset animatable for heads [3] and bodies [6, 53, 27] by rigging the 3D asset to a parameterized FLAME [35] or SMPL [41] mesh. Such generated 3D avatars are view consistent by design but lack the visual quality of 2D-based diffusion models. Furthermore, animation control is limited by the underlying mesh template. In our work, we employ a more accurate, implicit 3D morphable head model for better expression control and follow ControlNet [84] and IPAdapter [80] to build a diffusion-based neural renderer for high-quality image synthesis.
Figure 2: Method overview: We decode an NPHM expression code $z_{exp}$ in two ways to obtain a realistic image: We first extract an NPHM mesh and rasterize it from the desired viewpoint in (a), giving us canonical coordinates, depths, and normal renderings for the head mesh. In (b), the canonical coordinates $x_{can}$ are used to look up spatial features in a TriPlanes structure, rigging the features onto the mesh surface. Together with the rasterizer output, these mapped features form the input for the ControlNet part of DiffusionAvatar. The second route for the expression code goes through a linear layer depicted in (c). It yields expression tokens that are subsequently used in a newly added cross-attention layer inside the pre-trained latent diffusion model. Intuitively, the rasterized inputs should encode pose, shape and rough expression while the direct expression conditioning hints at more detailed facial expressions. The final image is synthesized in (d) by iteratively denoising Gaussian noise using the original DDPM denoising schedule [21].
3 Method
Our goal is to create a temporally consistent 3D head avatar of a person with explicit control over viewpoint and expression. We approach this task by designing a diffusion-based neural renderer that decodes learnable features rigged to the surface of an implicitly defined proxy geometry. In analogy to Deferred Neural Rendering [68], we dub this approach Deferred Diffusion. We begin by introducing the foundations of diffusion models (Sec. 3.1) and NPHM (Sec. 3.2) that we use as proxy geometry before going into detail about our method: In the rasterization (Sec. 3.3) and surface feature mapping (Sec. 3.4) stage, we generate the input images for our diffusion-based neural renderer. These renderings encode the desired viewpoint, head shape, and global expression. For more fine-grained expression control, we add NPHM’s expression codes as an additional input to the diffusion model (Sec. 3.5). Finally, our pipeline leverages a pre-trained diffusion model’s image synthesis and generalization capabilities for better quality.
3.1 Diffusion
We build our work upon latent diffusion models (LDM) [55], which operate in the the latent space of an autoencoder. Formally, LDMs employ an encoder $\mathcal{E}$ to map RGB images into a lower-dimensional latent space, facilitating generative tasks [12]. In the following, any mention of an image $x_{0}$ always refers to an RGB image $I$ that was mapped into the autoencoder’s latent space via $x_{0}=\mathcal{E}(I)$ to obtain a latent image.
In diffusion, the fixed forward process iteratively adds noise to the latent image:
$q(x_{\tau}|x_{\tau-1})$ $=\mathcal{N}\left(x_{\tau};\sqrt{\alpha_{\tau}}x_{\tau-1},(1-\alpha_{\tau})\mathbf{I}\right)$ (1)
where $\tau=1...N$ indicates the denoising step and the forward process variances $\alpha_{\tau}$ define the noise scheduling [21]. In practice, noisy samples $x_{\tau}$ are obtained with the standard Gaussian reparameterization:
$x_{\tau}=\sqrt{\bar{\alpha}_{\tau}}x_{0}+\sqrt{1-\bar{\alpha}_{\tau}}\epsilon\qquad\epsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)$ (2)
Typically, diffusion models are trained to predict the original noise $\epsilon$ given the noisy sample $x_{\tau}$. We empirically find a different method, known as the $v$-parameterization [57], to work better in our scenario, where $v$ is defined as:
$v$ $=\sqrt{\bar{\alpha}_{\tau}}\epsilon-\sqrt{1-\bar{\alpha}_{\tau}}x_{0}$ (3)
This has two advantages. First, $v$-prediction ensures that the loss can always guide the model to learn something meaningful, even when the input already contains a lot of noise, leading to faster convergence [22]. Second, it allows us to train the model also on pure noise inputs, which during inference is the most challenging denoising step. To this end, we rescale the noise schedule $\alpha$ to ensure zero signal-to-noise ratio inputs during training as proposed by [39].
3.2 NPHM
NPHM [17] is a morphable head model that generates a signed distance field (SDF) of a human head given an identity code $z_{id}$ and an expression code $z_{exp}$. We employ COLMAP [59, 58] to obtain pointclouds $\{\mathcal{P}_{t}\}$ for each timestep $t$ of a multi-view video sequence of the person. Subsequently, we fit NPHM to each pointcloud $\mathcal{P}_{t}$ to obtain $z_{id}$ and $z^{t}_{exp}$. We use a variant of NPHM that uses a backward deformation field instead of the originally proposed forward deformation field as it simplifies the fitting process. Formally, we obtain the NPHM fitting as follows:
$z_{id},z^{t}_{exp}=\operatorname*{arg\,min}_{z_{id},z_{exp}}\sum_{x\in\mathcal{P}_{t}}\left|\mathcal{F}_{id}(\mathcal{F}_{exp}(x))\right|$ (4)
where $\mathcal{F}_{id}$ is NPHM’s identity network implemented as a signed distance field, and $\mathcal{F}_{exp}$ is the backward deformation network. Note that we drop the dependency of $\mathcal{F}_{id}$ and $\mathcal{F}_{exp}$ on their respective latent codes for simplicity. The fitted SDF representation can then be translated into a mesh $M_{t}$ via marching cubes [42]. Furthermore, for each vertex $x\in M_{t}$ of the extracted mesh, we can obtain its canonical coordinates $x_{can}\in\mathbb{R}^{3+2}$ via the backward deformation field:
$x_{can}=\mathcal{F}_{exp}(x)$ (5)
where the first 3 coordinates of $x_{can}$ represent the usual spatial dimensions while the remaining 2 coordinates are ambient dimensions that can help resolve topological issues in mouth and eye regions [50]. We utilize the meshes ${M_{t}}$ in the subsequent rasterization stage, whereas the canonical coordinates $x_{can}$ are necessary for the spatial feature lookup.
3.3 Rasterization
We employ nvdiffrast [34] to generate the actual input images for the diffusion model. Let $M_{t}$ be a facial mesh of a person at timestep $t$ and $\pi_{t}$ a camera pose. We then obtain a set of renderings $R^{t}\in\mathbb{R}^{H\times W\times C}$ for this timestep using the rasterizer $\mathcal{R}$:
$R^{t}=\mathcal{R}\left(M_{t},\pi_{t}\right)$ (6)
In our case, the channels of $R^{t}$ are normals, depths, and a canonical coordinate rendering $R^{t}_{can}\in\mathbb{R}^{H\times W\times(3+2)}$ of the NPHM mesh.
3.4 TriPlane Feature Mapping
It has been shown that tying learnable features to a surface helps neural renderers to synthesize more detailed images, a technique known as neural textures [68]. Since NPHM lacks a consistent UV space due to its implicit nature, we propose a simple extension of neural textures: We tie learnable features to the surface of the extracted mesh $M_{t}$ by querying a spatial structure with the canonical coordinates $x_{can}$. We use TriPlanes [8] for the 3 spatial dimensions and a regular 2D feature map for the ambient dimensions:
$R^{t}_{\text{feat}}$ $=\textsc{TriPlane}\left(R^{t}_{can,0\text{-}3}\right)$ (7)
$R^{t}_{\text{feat\_amb}}$ $=\textsc{AmbientMap}\left(R^{t}_{can,3\text{-}5}\right)$ (8)
These learnable feature maps form the 2D input for our diffusion-based neural renderer, together with the other buffers $R^{t}$ from the rasterizer:
$R^{t}\leftarrow R^{t},R^{t}_{\text{feat}},R^{t}_{\text{feat\_amb}}$ (9)
3.5 Direct Expression Conditioning
Conditioning on renderings guides the head pose as well as coarse expressions. However, subtle details may not be easy for the network to decode from renderings. To facilitate the synthesis of detailed expressions, we add new cross-attention layers to the U-Net, following IPAdapter [80]. Let $Z$ be the intermediate feature map computed by an existing cross-attention operation in the pre-trained LDM: $Z=\textsc{attention}(Q,K,V)$. We then perform direct expression conditioning by adding another cross-attention layer:
$f^{t}_{exp}$ $=\textsc{exp}(z^{t}_{exp})$ (10)
$Z$ $\leftarrow Z+\textsc{attention}(Q,W^{k}f^{t}_{exp},W^{v}f^{t}_{exp})$ (11)
where exp is our expression conditioning module that linearly maps $z^{t}_{exp}$ into a sequence $f^{t}_{exp}\in\mathbb{R}^{4\times d}$ of 4 expression tokens forming the keys and values for cross-attention. In total, we add 15 cross-attention layers to the LDM.
3.6 Deferred Diffusion
The final rendering is obtained by iteratively denoising full noise $x_{T}\sim\mathcal{N}(0,I)$ with our diffusion-based neural renderer $\mathcal{D}$ conditioned on the rasterized NPHM meshes $R^{t}$ and expression tokens $f_{exp}$:
$x_{\tau-1}\sim\mathcal{D}(x_{\tau})=\mathcal{S}(x_{\tau},f^{t}_{exp},\mathcal{C}(x_{\tau},R^{t}))$ (12)
where $\mathcal{C}$ is the ControlNet architecture and $\mathcal{S}$ a pre-trained LDM with our expression conditioning modules inserted. The denoised prediction $x_{\tau-1}$ is obtained via the sampling procedure of [21].
During training, we minimize the following loss:
$\mathcal{L}=\mathbb{E}_{\epsilon,\tau,x^{t}_{0},R^{t},f^{t}_{exp}}\left[\left\|\mathcal{D}\left(x^{t}_{\tau},R^{t},f^{t}_{exp}\right)-v\right\|_{2}\right]$ (13)
where $\epsilon$ is the sampled noise, $\tau$ is the denoising step and $x^{t}_{0}=\mathcal{E}(I^{t})$ is the ground truth latent image. We use $\mathcal{L}$ to optimize (i) the ControlNet $\mathcal{C}$, (ii) the parameters of expression conditioning $W^{k},W^{v}$, and exp, and (iii) the spatial feature maps TriPlane and AmbientMap.
NeRFace [15]
DNR [68]
DNR + GAN
DiffusionRig [11]
Ours
GT
Figure 3: Qualitative results for self-reenactment. We compare against 3D methods (NeRFace [15]) and methods employing 2D renderers (Deferred Neural Rendering [68], DiffusionRig [11]). Note that the slightly gray background for DiffusionRig is caused by their training scheme using $\epsilon$-prediction (see Sec. 3.1). Our method consistently produces more expressive facial performances while simultaneously providing more detailed renderings.
Source Actor
NeRFace [15]
DNR [68]
DNR + GAN
DiffusionRig [11]
Ours
Figure 4: Qualitative results for avatar animation. Our method faithfully transfers the source actor’s expression and consistently produces compelling renderings, even for complex performances.
4 Experimental Results
4.1 Training
For our experiments, we employ the recently released multi-view video dataset from NeRSemble [33]. The dataset contains 26 different multi-view sequences per person captured with 16 cameras, featuring different facial expressions, sentences and emotions. We use BackgroundMattingV2 [38] to remove the background and a segmentation network to remove the torso [86]. We fit NPHM on the video sequences of 5 different persons to create pairings of multi-view videos and tracked meshes. In total, we use about 3300 timesteps per person. We train one DiffusionAvatar for each person on these pairings with Adam [32] using a learning rate of $41e-$ 4 for ControlNet and the expression conditioning layers, and a learning rate of $21e-$ 2 for the TriPlane feature lookups. We use a batch size of $B=8$ and spatial resolutions $H=W=512$. We employ random cropping and resizing of the target images to avoid overfitting. The size of our spatial feature maps is $512\times 512\times 16$ for each TriPlane and the ambient map. In total, this yields 73 channels for the input renderings $R_{t}$ fed to our diffusion-based renderer. The pre-trained LDM is a Stable Diffusion v2.1 network [55]. Training takes roughly two days on a single RTX A6000 for 100k gradient descent steps.
4.2 Experiment Setup
Tasks.
We measure the quality of a 3D avatar in terms of image quality, 3D consistency and expression synthesis. To this end, we conduct experiments on two tasks: Self-reenactment and avatar animation. For self-reenactment, we hold out 2 complete sequences and the frontal camera of every sequence during training to measure both image quality and view consistency. During testing, we provide the holdout tracked meshes of the same person and render from the novel view. For avatar animation, we first generate NPHM meshes using the same identity code $z_{id}$ as the avatar and a sequence of expression codes $\{z^{t}_{exp}\}$ from a different person. We then render a video of the avatar from a novel view. To evaluate the realism of this expression transfer, we conduct a user study where we show the animated avatars to participants.
Metrics.
We employ two paired-image metrics to measure the quality of individual generated frames: Peak Signal-to-Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [85]. Furthermore, to assess temporal consistency, we employ the perceptual video metric JOD [43] to measure visual similarity of a rendered video to its ground truth counterpart. Finally, we make use of two face-specific metrics: Average Keypoint Distance (AKD) measured in pixels with keypoints estimated from PIPNet [28], and cosine similarity (CSIM) of identity embeddings based on ArcFace [9].
Baselines.
We compare our method against the following state-of-the-art 3D head avatar creation methods:
NeRFace [15]. A NeRF-based method that reconstructs a radiance field conditioned on expression codes. The original implementation is for monocular videos and uses FLAME. We extend it to our multi-view training scenario and provide NPHM expression codes for better comparison.
Deferred Neural Rendering (DNR) [68]. A screen-space decoder built on top of Pix2Pix [89] that uses learnable neural textures rigged to a tracked Basel Face Model [51]. We provide our tracked NPHM meshes as input for better comparison. Additionally, we compare against a version of DNR that also uses an adversarial loss for sharper images that we call DNR+GAN in the following.
DiffusionRig [11]. A recent diffusion-based method that uses albedo, normal, and shaded color renderings of FLAME meshes tracked by DECA [13] as input to control pose and expression. While their method was originally proposed to be fine-tuned on a small personal photo album of  20 images, we fine-tune their pre-trained model on all training frames of our dataset.
4.3 Self-Reenactment
In Tab. 1, we compare DiffusionAvatars to other state-of-the-art approaches on the self-reenactment task. Results are averaged over avatars generated from five different persons. Our method shows the overall best performance with the biggest improvement in perceptual and face-specific metrics. The results indicate that DiffusionAvatars generates sharper images (LPIPS) that are more temporally consistent (JOD). Furthermore, our method better maintains the person’s identity when rendered from a novel view (CSIM) than the baselines. Finally, facial features such as the position of mouth corners and eyebrows synthesized by our method match the ground truth expressions more closely (AKD).
We further provide a qualitative comparison in Fig. 3. The baselines struggle especially in facial regions that are poorly explained by the NPHM mesh, such as the mouth interior. In contrast, our method can plausibly fill these regions, providing a consistently realistic appearance.
Method Self Reenactment User Study
[…]
NeRFace [15] 22.5 0.299 6.65 5.10 0.761 2.19 2.35
DiffusionRig [11] 19.6 0.218 6.42 2.66 0.886 2.47 3.52
DNR [68] 24.4 0.226 7.31 2.18 0.898 2.87 3.94
DNR+GAN [68] 22.8 0.115 7.06 2.32 0.859 3.06 3.97
Ours 24.5 0.088 7.53 1.86 0.912 4.02 4.14
Table 1: Quantitative Comparison: We report metrics for a self-reenactment scenario on unseen expressions and unseen views. For our user study on avatar animation, we collect 735 responses from 49 participants to measure Visual Quality (VQ) and Driving Fidelity (DF) on a scale of 1-5.
4.4 Avatar Animation
Due to the task’s inherent lack of any ground truth, we perform a user study to validate avatar animation. We collect 735 responses from 49 participants on two questions: (i) Visual Quality (VQ): What is the overall quality of the avatar animation? (ii) Driving Fidelity (DF): How closely does the animated avatar match the expressions of the driving sequence? We present two videos side-by-side: One sequence of a source actor performing various expressions, and a video of another person’s avatar with the expressions transferred to it. The results of our user study can be seen in Tab. 1. Our method achieves the highest scores for both visual quality and similarity to the driving sequence. The difference is particularly pronounced for the visual quality score. We attribute this to the fact that users heavily penalize any visual artifacts immediately obvious in video renderings.
These perceptual differences in the task of avatar animation are also apparent in the qualitative comparison shown in Fig. 4.
4.5 Ablations
In the following, we examine the design choices of our method. All ablation experiments are conducted on 3 avatars on the self-reenactment task, and the averaged results are summarized in Tab. 2.
[…]
w/o diffusion 25.1 0.133 7.69 1.96 0.892
FLAME 24.2 0.083 7.36 2.46 0.900
w/o exp. cond. 24.5 0.081 7.65 1.98 0.911
w/o LDM prior 24.5 0.078 7.67 1.89 0.913
w/o spatial features 24.9 0.078 7.77 1.87 0.918
spherical UV 24.9 0.075 7.77 1.95 0.920
Ours 25.3 0.074 7.85 1.91 0.918
Table 2: Ablation of architectural choices. Our full method utilizing TriPlane feature mapping performs overall the best. The biggest quantitative improvements stem from using NPHM instead of FLAME, employing direct expression conditioning, and using a pre-trained LDM.
Effect of Diffusion.
We study whether diffusion models are necessary for our task. To this end, we train our architecture without the pre-trained LDM to directly predict the output in a single forward pass, similar to Deferred Neural Rendering. The results indicate that such a method performs notably worse, especially in the LPIPS metric, indicating that diffusion is important to obtain sharp renderings.
Effect of NPHM vs FLAME.
We compare DiffusionAvatars with a version trained on renderings of a fitted FLAME instead of NPHM. Since NPHM models facial geometry as an SDF, it possesses a much higher representational power than FLAME. As a result, the renderings are much closer to the actual desired appearance of the head. This is also reflected by the results of our ablation study, where using FLAME gives consistently worse performance. This is especially true for the AKD metric, indicating that our diffusion-based neural renderer cannot reenact certain complex expressions based on FLAME alone.
Effect of Expression Conditioning.
The overall performance worsens when direct expression conditioning is turned off. We attribute this to the fact that expression conditioning helps to distinguish subtle expression details that can hardly be inferred from the NPHM mesh alone. It also helps DiffusionAvatars to synthesize expressions not captured by NPHM, such as tongue movement, by enabling the network to correlate such appearances with certain expression code combinations.
Effect of 2D LDM Prior.
We train our pipeline from scratch without resorting to a pre-trained LDM. In this case, we remove the ControlNet module and condition the U-Net directly on the NPHM renderings. This architecture, trained from scratch, performs competitively but slightly worse. We see the biggest advantage of pre-training in faster convergence and better coherence of the generated images.
Effect of Spatial Features.
We compare three spatial feature mapping approaches: (i) Our Triplanes, (ii) a simpler approach based on spherical feature mapping, and (iii) not using any spatial features at all. For spherical feature mapping, we create an approximate UV map for each NPHM mesh by projecting the 3D canonical coordinates for each vertex onto a sphere centered inside the head. Tab. 2 shows that using TriPlanes boosts most metrics, whereas the simpler spherical mapping brings almost no benefit.
5 Limitations
DiffusionAvatars can create photo-realistic 3D head avatars with pose and expression control. More challenges must be solved, however, before an avatar can be used in production. For example, to put an avatar into a realistic environment, it is necessary to have control over the lighting properties. Currently, DiffusionAvatars bakes the lighting into the generated images. Since we employ a detailed 3D head geometry underneath, one could imagine directly modeling shadow effects as a composite of the synthesized images. Further, our current architecture is not (yet) amenable to real-time applications due to the comparatively compute-intensive denoising loop. Here, recent advances in the distillation of diffusion models may provide a remedy [61, 44].
6 Conclusion
This paper presents a novel method DiffusionAvatars, which creates a photo-realistic 3D head avatar from multi-view videos using a tracked parametric head model of a person. DiffusionAvatars can be animated by a source actor or directly via NPHM. We use rendered NPHM meshes with rigged spatial features as input to a diffusion-based neural renderer. We leverage the image synthesis capabilities of a pre-trained latent diffusion model with ControlNet, facilitating generalization to unseen expressions. While this pipeline is already capable of synthesizing appealing images, we show that additionally conditioning the neural renderer on NPHM’s expression codes further improves the model’s ability to generate complex facial expressions. Our experiments suggest that the proposed architecture can effectively synthesize high-quality renderings and animate the 3D avatar with high fidelity. We believe our approach demonstrates an exciting application for high-fidelity 3D avatar creation.
Acknowledgements
This work was supported by the ERC Starting Grant Scan2CAD (804724) and the German Research Foundation (DFG) Research Unit “Learning and Simulation in Visual Computing”. We would also like to thank Karla Weighart for proof-reading and Angela Dai for the video voice-over.
Appendix
Appendix A Novel View Synthesis Results
[…]
NeRFace [15] 28.12 0.284 7.93 2.72 0.841
DiffusionRig [11] 19.73 0.231 6.02 9.71 0.917
DNR [68] 26.90 0.213 7.91 1.58 0.933
DNR+GAN [68] 23.56 0.104 7.30 1.70 0.906
Ours 26.93 0.067 8.28 1.16 0.939
Table 3: Quantitative Comparison: We report metrics for novel-view-synthesis.
Tab. 3 contains results for novel view synthesis. We measure the performance from the unseen frontal view and average over the 24 train sequences for each of our 5 avatars. DiffusionAvatars scores the best metrics except for PSNR, which favors the overly blurry renderings of NeRFace. These results show that our method also provides good 3D control.
Appendix B User Study Setup
In  Fig. 5, we show the interface of our user study. We ask users to rate visual quality and driving fidelity by providing them with the following definitions:
Visual Quality.
The overall quality of the generated video: Is there any flickering, visual artifacts, blurriness, etc., that feel unnatural?
Driving Fidelity.
How closely does the generated video follow the desired facial expressions: Are the displayed expressions the same? Is the emotion displayed by the avatar recognizable?
During the study, each user is presented with 15 randomly chosen pairs of driving sequences (left) and avatar animations (right). The pairs were drawn from 135 different avatar animations generated by the baselines and our method. In total, we received 735 responses for each question from 49 participants.
Figure 5: Setup of our User Study. We ask participants to evaluate the avatar animation task in which an avatar is controlled by a sequence of a different person.
Appendix C Qualitative Comparison of Ablations
As an addition to Tab. 2 of the main paper, we provide qualitative comparisons of three ablations in Fig. 6, Fig. 7 and Fig. 8.
Appendix D Societal Impact
DiffusionAvatars provides the means to generate realistic images of faces with control over viewpoint, pose, and expression. As such, there is a risk of misusing an avatar to the original person’s disadvantage. However, we focus on 3D applications such as immersive teleconferencing or character animation, not indistinguishable 2D video synthesis. Our method does not model lighting, the background, or the torso. Therefore, we believe a detection system can be built to identify generations of our method by exploiting such characteristics [56].
All subjects shown in the paper and the supplemental video are part of the NeRSemble dataset [33] and have consented to their recordings being used for research purposes.
FLAME [35]
Ours
GT
Figure 6: Effect of FLAME vs. NPHM. Using FLAME instead of NPHM as the underlying morphable head model makes the generated faces less expressive. This is due to FLAME’s limited geometry and expression space, which does not provide enough geometric cues for our diffusion-based neural renderer.
w/o exp. cond.
Ours
GT
Figure 7: Effect of expression conditioning. Directly providing $z_{exp}$ to the diffusion model helps synthesize areas not modeled by NPHM’s geometry, such as eye movement.
w/o LDM prior
Ours
GT
Figure 8: Effect of LDM prior. Using a pre-trained LDM instead of training the diffusion model from scratch steers the generation towards more realistic faces.